OpenStudy (anonymous):

Find the equation of the line tangent to the graph of f at (3, 140), where f is given by f(x) = 8x^3 − 9x^2 + 5.

OpenStudy (aravindg):

just differentiate the cubic equation to get f'(x)

OpenStudy (aravindg):

now f'(3) will give you slope of tangent

OpenStudy (anonymous):

24x^2-18x

OpenStudy (aravindg):

now you have slope of tangent and a point in the tangent , use slope point form to write equation of tangent

OpenStudy (anonymous):

f'(3)?

OpenStudy (aravindg):

yep substite x=3 in the differential

OpenStudy (anonymous):

so slope is 162

OpenStudy (anonymous):

so now the mx+b?

OpenStudy (aravindg):

slope point form y-y1=m(x-x1)

OpenStudy (aravindg):

gt it?

OpenStudy (anonymous):

yeah

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

y-140=162(x-3)?

OpenStudy (anonymous):

right

OpenStudy (aravindg):

yep

OpenStudy (anonymous):

thanks

OpenStudy (aravindg):

welcomes